# Navigating the Landscape of Hint Generation Research: From the Past to the Future

### v1

@@ 1 Introduction
Prior research has established a correlation between the student-teacher ratio and student’s overall performance Koc and Celik (2015). However, private tutoring is not accessible to everyone, and finding expert tutors is often difficult and incurs considerable costs Bray (1999); Graesser et al. (2012). Intelligent tutoring systems (ITSs) holds the key to addressing these educational challenges, notably the need for personalized learning in a system often reliant on instructional teaching and standardized testing Anderson et al. (1985).
@@
Within education, hints refer to the clues, prompts, questions, or suggestions provided to learners to aid them in solving problems, answering questions, or completing tasks, thereby encouraging critical thinking, problem-solving skills, and independent learning . In Figure 1, we provide an example of a hint generation system capable of the reasoning required to answer the question, acknowledging the wrong attempt by the learner, and providing informative hint linked to learner’s existing knowledge. This framework of scaffolding is well-established in education Van de Pol et al. (2010), and expert tutors are guided to incorporate it in their teaching practices Belland (2017).
With the aim of developing hint generation systems with capabilities such as the ones showcased in Figure 1, we consolidate the dispersed efforts on hint generation bridging the gap between research in education and cognitive sciences, and research in AI and Natural Language Processing (NLP). Grounded in the findings from our literature review, we provide a roadmap for future research in hint generation. We summarize the key characteristics of a successful hint as observed by research with human tutors in Section 2 and review the automated hint generation systems in Section 3. We identify the gaps and propose a roadmap for future research in hint generation in Section 4. We provide a rethinking of the formal definition of the hint generation task (Section 4.1), a brief review of research areas that can inform the design of a hint generation system that aligns with the formal definition (Section 4.2), and open challenges for future directions for effective hint generation systems (Section 4.3) and last but not least ethical considerations (Section 4.4) . Our major contributions include:
@@
    A literature review on hint generation that bridge the gap between research in education and cognitive science and research in AI and NLP.
@@
    A roadmap for research on hint generation, outlining challenges, promising future directions and ethical considerations for the field.
@@ 2.3 Style of a hint
Expert human tutors adopt diverse techniques to convey information to learners. These strategies vary from non-verbal cues such as body language, facial expressions, and vocal tone Bambaeeroo and Shokrpour (2017); Wahyuni (2018) to adopting multimedia content to teach complex topics (e.g. using animations and maps to teach the geographical concept of folded mountains) Kapi et al. (2017). We try to cover the most relevant aesthetic aspects of hints that might be useful in building better hint-generation system.
@@ 3 Survey of Computational Approaches
In this section, we provide a comprehensive overview of the recent advancements of computational approaches for automatic hint generation. We first describe the extensively studied hint generation techniques for computer programming (learning how to program), and then dive into question answering-based hint generation, where we explore natural language hint generation strategies for diverse domains like mathematics, language acquisition, or factual entity-based questions. We conclude this section by describing some limitations of hint generation systems today, and potential directions for improvement.
@@ 3.1 Hint Generation for Computer Programming
A vast majority of computational approaches for hint generation have been focused on the specific domain of computer programming, owing to the objective nature of the task and abundance of data. We briefly discuss the approaches, datasets, and evaluation metrics adopted in the field. For a more comprehensive review in this specific domain, we refer the readers to the surveys written by Le et al. (2013), Crow et al. (2018), McBroom et al. (2021) and Mahdaoui et al. (2022).
Datasets. Two widely popular datasets in the programming hint generation space are iSnap Price et al. (2017) and ITAP Rivers and Koedinger (2017) datasets. Both datasets consist of detailed logs collected from several students working on multiple programming tasks, including the complete traces of the code and records of when the hints were requested. iSnap Price et al. (2017) is based on the Snap! [^1] https://snap.berkeley.edu/ – a block-based educational graphical programming language, while ITAP Rivers and Koedinger (2017) is a Python dataset collected from two introductory programming courses taught at Carnegie Mellon University. In Table 1, we describe an example from the ITAP dataset, where the goal is to write a program to determine if a given day is weekend. Given a student’s program that fails to pass the pre-determined test cases (e.g., isWeekend(“Sunday") will return False), the aim of a hint generation system is to provide hints to help them successfully solve the problem (e.g., replacing lowercase ‘saturday’ to uppercase ‘Saturday’).
@@
Approaches. Most of the recent efforts in programming hint generation adopt a data-driven deterministic approach Barnes and Stamper (2008); Rivers and Koedinger (2017); Obermüller et al. (2021); Jin et al. (2012); Zimmerman and Rupakheti (2015); Paaßen et al. (2018); Price et al. (2016); Rolim et al. (2017), that comprises of three key components - a corpus of diverse candidate solutions (usually obtained via past student attempts), a matching algorithm to select the best candidate response given an ongoing attempt based on similarity, and graph-based solution path construction to synthesize hints (Fig 2 describes one approach in detail). We found abstract syntax tree (ASTs) McCarthy (1964); Knuth (1968) to be the most popular choice of graph-representation for hint synthesis, due to its vast literature and language-agnostic nature. Similarly, McBroom et al. (2021) provide a detailed generalization of hint generation techniques for the programming domain called HINTS (Hint Iteration by Narrow-down and Transformation Steps framework). Although effective, the programming hints are rarely natural language responses, and therefore are not capable of incorporating the stylistic aspects of hints that enhance the learner’s experience (Section 2.3). We believe an amalgamation of NLP technologies and current hint generation systems could improve upon this limitation.
Evaluation metrics. Price et al. (2019) develop QualityScore, an evaluation framework that quantifies the quality of programming hints by comparing the generated hints against expert tutor written hints using abstract syntax tree McCarthy (1964); Knuth (1968) based similarity overlap measure. They benchmark the most prominent data-driven techniques on the iSnap Price et al. (2017) and ITAP Rivers and Koedinger (2017) datasets, and found the evaluation framework to align well with the expert opinions. Price et al. (2019) also identified a few limitations of the benchmarked hint generation systems, such as, a) difficulty identifying the focus of hint request in large code, b) inability to handle divergent code that does not match the existing corpus of past student attempts, and c) inability to understand the student intent.
@@
3.2 Question Answering based Hint Generation
Question answering format is the most prevalent setting for hint generation, where the learner is trying to answer a question to recall and concertize their understanding about a concept. Recent years have seen an increased effort in hint generation frameworks beyond programming hint generation.
@@ 3.2 Question Answering based Hint Generation
Datasets. ReMath Wang et al. (2023) is a benchmark developed for evaluating and tutoring students in the mathematics domain. ReMath provides a systematic breakdown of the human-tutoring process into three steps - 1) identifying the error type, 2) determine a response strategy and intention, and 3) generate a feedback response that adheres to this strategy (example in Table 1). Wang et al. (2023) also provide a set of error types (e.g., guess, careless, misinterpret, right-idea), response strategies (e.g., explain a concept, ask a question) and intentions (e.g., motivate the student, get the student to elaborate the answer) to facilitate the feedback generation process. ReMath is a great example of high-quality data collection with human experts towards building better hint generation systems.
@@
Approaches. For open-ended hint generation of factoid questions, Jatowt et al. (2023) proposed a Wikipedia-based retrieval framework for the “Who?”, “Where?” and “When?” type questions. They propose a popularity-based framework for the "When" question type, where the popularity of an event for the answer year is measured by the count of Wikipedia hyperlinks directing to the event’s website, and a hand-curated template approach for the "Who" and "Where" question types (example hint described in Table 1). On the other hand, Wang et al. (2023) benchmarked the ReMath dataset by instruction fine-tuning Wei et al. (2021) the language models like Flan-T5 Chung et al. (2022) and GODEL Peng et al. (2022) and using in-context learning Dong et al. (2022) prompts for gpt-3.5-turbo and gpt-4 Achiam et al. (2023). Both Tack and Piech (2022) and Wang et al. (2023) found the direct use of large language models to fall short in comparison to human expert responses. Current question-answering based hint generation systems don’t personalize the hints to learner’s preferences, learning objectives, or consider their prior knowledge (Sections 2.1 and 2.2). We discuss how we can improve these hint generation systems going forward in Section 4.
@@
Evaluation Metrics. All the discussed approaches for hint generation have use human evaluation to assess the quality of the system’s output. Jatowt et al. (2023) conducted a between subjects study evaluating the baselines and proposed hint generation strategy across different groups. Tack and Piech (2022) proposed the "AI Teacher Test", comparing the generated responses against the teacher responses across three dimensions - “speak like a teacher”, “understand a student”, and “help a student”. They identified that the LLMs are good at conversation uptake (i.e., the first two requirements), but are quantifiably worse than real teachers on several pedagogical dimensions, especially helpfulness to a student. Wang et al. (2023) evaluate the error type identification and feedback response strategy selection as a multi-class classification task, and utilize exact match and Cohen’s kappa to measure the accuracy, and entropy to measure the output diversity. They also conduct human evaluations for the response generation task, and found all models constrained with knowledge of ground-truth error type and response strategy outperform their unconstrained counterparts.
@@ 4 Roadmap for Future Research in Hint Generation
Great progress has been made in hint generation over the last two decades; however, there is still room for improvement. The hint generation systems today do not personalize the hints to the learner’s prior knowledge (Section 2.1), and are only studied in short-term studies. We still do not know the effects of long-term exposure to these systems on the learners. These frameworks are also limited to certain domains, and have not been much explored in other domains including different branches of science and social science. To improve upon these factors, we discuss a roadmap for the future efforts in hint generation. We propose a hint generation framework that draws on education and cognitive sciences (Figure 3).
Figure 3: Roadmap of the proposed hint generation system.
@@ 4.1 Formal Definition
Jatowt et al. (2023) propose a formal definition for hint generation as follows: Given a question $q$ and its correct answer $a$, the task is to generate a hint $h$, such that $P(a|q,h)-P(a|q)>\epsilon$, where $P(a|q,h)$ denotes the probability of a user answering the $q$ if the hint $h$ is given, $P(a|q)$ is the probability of a user answering the $q$ without the hint, and $\epsilon$ is a threshold parameter ($\epsilon>0$).
This definition only emphasizes the hint’s ability to help answer a question, and does not incorporate any pedagogical aspects or principles. It also does not take the individual preferences into considerations. Below, we provide a more generic definition, that draws inspiration from the widely adopted cognitive frameworks on human learning such as Anderson’s adaptive control of thought rational (ACT-R) theory Anderson (2013) and Ausubel’s theory on meaningful learning Ausubel (1963, 1962, 2012). We also incorporate the findings from the qualitative experiments discussed in Section 2, explicitly integrating the alignment of hints to student’s learning objective, and their prior knowledge (Section 2.2), as well as incorporating the pragmatics of a hint (Section 2.1) by accounting for learner’s preferences. We formulate the hint generation task in a tutoring framework, with the aim to correctly answer a question. We will later explain how the definition can be extended to generating hints for tasks beyond question answering.
@@
Given, a learner $l$ attempting to answer a question $q$, a hint generation system H (Figure 3) generates a hint $h\in\mathcal{H}$ by mapping $\textbf{H}:\mathcal{I}\rightarrow\mathcal{H}$, where =$\mathcal{I}=\{q,a,\mathcal{K}_{q\rightarrow a},D_{q}^{l},\mathcal{L}_{l},\mathcal{F}^{l}_{learning},\mathcal{F}^{l}_{pref}\}$ is the input to the hint generation system having the following elements:
  • $q$: the question
  • $a$: the correct answer
@@
  • $D_{q}^{l}=\{q,\hat{a}_{1},\hat{h}_{1},\hat{a}_{2},\hat{h}_{2},...,\hat{a}_{k},\hat{h}_{k}\}$: is the ongoing dialogue, where $\hat{a}_{i}$ and $\hat{h}_{i}$ are, respectively, the learner’s past attempts and hints related to $q$,
  • $\mathcal{L}_{l}=\{D_{q_{i}}\}$: the learner $l$’s past learning history,
@@
Such that the generated hint h doesn’t contain the answer (Equation 1), helps the learner to answer the question (Equation 2), and aligns with the learner’s learning objective(s) (Equation 3).
@@
Both Anderson’s ACT-R theory and Ausubel’s meaningful learning theory imply the significance of connecting the new knowledge required to solve a problem to the existing concepts in learner’s knowledge base. We define the hint generation task (Equation 4) to incorporate these findings. We also account for the diversity of learners’ motivation to study, and incorporate the notion of improving an individual’s learning objective (Equation 3). The hint generation strategy for a learner aiming to improve their answer accuracy would greatly differ from someone aiming to maximize the diversity of acquired knowledge in a learning session.
Extending the formal definition. The given definition although generic, assumes a question-answering setup for the hint generation task, with an additional assumption of the availability of objective answers. We can modify this definition to accommodate for other hint generation settings as described below.
@@
  • For a hint generation system for the writing assistance task, we need to replace the question with the writing task description $q\rightarrow T$ and ongoing dialogue with an interactive sequence of learner’s writing $\hat{a_{i}}\rightarrow\hat{w_{i}}$, and past hints $\hat{h_{i}}$ $\mathcal{D}_{q}^{l}\rightarrow\mathcal{W}_{T}^{l}\mid\mathcal{W}_{T}^{l}\Leftrightarrow\{\hat{w_{1}},\hat{h_{1}},\hat{w_{2}},\hat{h_{2}},...\}$, and the answer with a target rubric for the writing evaluation $a\rightarrow\mathcal{R}$ to get the formal definition of a hint generation system for the writing assistance task.
@@
4.2 Components of an Effective Hint Generation System

### v2

@@ 1 Introduction
Prior research has established a correlation between the student-teacher ratio and student’s overall performance Koc and Celik (2015). However, private tutoring is not accessible to everyone, and finding expert tutors is often difficult and incurs considerable costs Bray (1999); Graesser et al. (2012). Intelligent tutoring systems (ITSs) hold the key to addressing these educational challenges, notably the need for personalized learning in a system often reliant on instructional teaching and standardized testing Anderson et al. (1985).
@@
Within learning science, hints refer to the clues, prompts, questions, or suggestions provided to learners to aid them in solving problems, answering questions, or completing tasks, thereby encouraging critical thinking, problem-solving skills, and independent learning. In Figure 1, we provide an example of a hint generation system capable of the reasoning required to answer the question, acknowledging the wrong attempt by the learner and providing informative hints linked to the learner’s existing knowledge. This framework of scaffolding is well-established in education Van de Pol et al. (2010), and expert tutors are guided to incorporate it in their teaching practices Belland (2017).
With the aim of developing hint generation systems with capabilities such as the ones showcased in Figure 1, we consolidate the dispersed efforts on hint generation, bridging the gap between research in education and cognitive sciences on the one hand, and research in AI and Natural Language Processing (NLP) on the other. Grounded in the findings from our literature review, we provide a roadmap for future research on automatic hint generation. We summarize the key characteristics of a successful hint as observed by research with human tutors in Section 2 and review the automated hint generation systems in Section 3. We identify the gaps and propose a roadmap for future research in hint generation in Section 4. We provide a rethinking of the formal definition of the hint generation task (Section 4.1), a brief review of research areas that can inform the design of a hint generation system that aligns with the formal definition (Section 4.2), open challenges for future directions for effective automatic hint generation systems (Section 4.3), and ethical considerations (Section 4.4). Our major contributions include:
@@
    A literature review on hint generation that bridges the gap between research in education and cognitive science on the one hand and research in AI and NLP on the other.
@@
    A roadmap for research on automatic hint generation, outlining challenges, promising future directions, and ethical considerations for the field.
@@ 2.3 Style of a hint
Expert human tutors adopt diverse techniques to convey information to learners. These strategies vary from non-verbal cues such as body language, facial expressions, and vocal tone Bambaeeroo and Shokrpour (2017); Wahyuni (2018) to adopting multimedia content to teach complex topics (e.g. using animations and maps to teach the geographical concept of "folded mountains") Kapi et al. (2017). We try to cover the most relevant aesthetic aspects of hints that might be useful in building better hint-generation system.
@@ 3 Survey of Computational Approaches
In this section, we provide a comprehensive overview of the recent advancements of computational approaches for automatic hint generation. We first describe the extensively studied hint generation techniques for computer programming that focus on revealing code snippets to help learn how to program. Next, we dive into the relatively under-explored natural language hint generation, where we explore strategies for diverse domains like mathematics, language acquisition, or factual entity-based questions. We conclude this section by describing some limitations of automatic hint generation systems today, and propose a roadmap for future research in the field in Section 4.
@@ 3.1 Hint Generation for Computer Programming
A vast majority of computational approaches for hint generation have focused on the specific domain of computer programming, owing to the more objective nature of the task and abundance of data. We briefly discuss the approaches, datasets, and evaluation metrics adopted in the field. For a more comprehensive review in this specific domain, we refer the readers to the surveys written by Le et al. (2013), Crow et al. (2018), McBroom et al. (2021) and Mahdaoui et al. (2022).
Datasets. Two widely popular datasets in the programming hint generation space are iSnap Price et al. (2017) and ITAP Rivers and Koedinger (2017) datasets. Both datasets consist of detailed logs collected from several students working on multiple programming tasks, including the complete traces of the code and records of when the hints were requested. iSnap Price et al. (2017) is based on Snap! [^1] https://snap.berkeley.edu/ – a block-based educational graphical programming language, while ITAP Rivers and Koedinger (2017) is a Python dataset collected from two introductory programming courses taught at Carnegie Mellon University. In Table 1, we describe an example from the ITAP dataset, where the goal is to write a program to determine if a given day is weekend. Given a student’s code that fails to pass the pre-determined test cases (e.g., isWeekend(“Sunday") will return False), the aim of a hint generation system is to provide hints to help them successfully solve the problem (e.g., replacing lowercase ‘saturday’ to uppercase ‘Saturday’).
@@
Approaches. Most of the recent efforts in programming hint generation adopt a data-driven deterministic approach Barnes and Stamper (2008); Rivers and Koedinger (2017); Obermüller et al. (2021); Jin et al. (2012); Zimmerman and Rupakheti (2015); Paaßen et al. (2018); Price et al. (2016); Rolim et al. (2017), that comprises of three key components: a corpus of diverse candidate solutions (usually obtained via past student attempts), a matching algorithm to select the best candidate response given an ongoing attempt based on similarity, and graph-based solution path construction to synthesize hints (Fig 2 describes one approach in detail). We found abstract syntax tree (ASTs) McCarthy (1964); Knuth (1968) to be the most popular choice of graph representation for hint synthesis due to its vast literature and language-agnostic nature. Similarly, McBroom et al. (2021) provide a detailed generalization of hint generation techniques for the programming domain called HINTS (Hint Iteration by Narrow-down and Transformation Steps framework). Although effective, programming hints are rarely natural language responses and, therefore, are not capable of incorporating the stylistic aspects of hints that improve the learner’s experience (Section 2.3). We believe an amalgamation of NLP technologies and current hint generation systems could improve upon this limitation.
Evaluation metrics. Given the vast majority of work adopting a data-driven hint strategy, programming hints are designed to reveal some aspect of the program not written by the learner. Within our survey, all the evaluation metrics adopted a code-based similarity measure to gauge the quality of the hint, and we found two prominent evaluation paradigms that we categorize into reference-based and reference-free evaluation metrics. The reference-based evaluation metrics Price et al. (2019) assume the availability of reference hints developed by expert tutors. For instance, QualityScore Price et al. (2019) is a reference-based evaluation metric that uses abstract syntax tree-based similarity measure to evaluate the quality of the generated programming hints with respect to expert tutor written hints.
Reference-free evaluation metrics Rivers and Koedinger (2017); Paaßen et al. (2018); Obermüller et al. (2021); Zimmerman and Rupakheti (2015) focus on measuring the impact of a hint to the learner’s response. For example, Paaßen et al. (2018) proposed root-mean-square error over two distance measures: (1) distance between the predicted post-hint state and true next state, and (2) distance between the predicted post-hint state and learner’s true final state. Rivers and Koedinger (2017), on the other hand, adopted a learner-agnostic approach to evaluate the hint quality. They used two measures of a successful chain of hints: (1) the ability of the hint sequence to reach the correct solution state and (2) the length of the chain.
Reference-based evaluation metrics are able to compare the quality of generated hints with the expert tutor’s feedback capabilities, but can be difficult to scale. Reference-free evaluation metrics, on the other hand, are capable of evaluating previously unseen problems but require human evaluations to obtain performance signals for the metrics. Each of these paradigms has pros and cons, but currently, there is no holistic measure of the quality of programming hints that emphasizes the pragmatics, semantics and style characteristics of the generated hints.
@@
TrivialQA
Question-
In which city are the headquarters of the International Monetary Fund?
Expected Answer-
Washington D.C.
Hint Generated by Mozafari et al. (2024b)-
The city is known for its neoclassical architecture.
The city is located on the Potomac River.
The city is the capital of the USA located on the east coast.
@@
3.2 Natural Language Hint Generation
Natural language hint generation has risen to popularity in the last few years as a consequence of the recent advancements in NLP, particularly large language models capable to generate fluent and coherent text Min et al. (2023); Naveed et al. (2023). We found the question answering format to be the most prevalent setup for natural language hint generation systems, where the learner attempts to answer a question to recall and concertize their understanding of a concept.
@@ 3.2 Natural Language Hint Generation
Datasets. ReMath Wang et al. (2023) is a benchmark co-developed with math teachers (i.e., experts) for evaluating and tutoring students in the mathematics domain. ReMath provides a systematic breakdown of the human-tutoring process into three steps: 1) identifying the error type, 2) determining a response strategy and intention, and 3) generating a feedback response that adheres to this strategy (example in Table 1). Each of these steps is manually annotated by an expert math teacher. Wang et al. (2023) also provides a set of error types (e.g., guess, careless, misinterpret, right-idea), response strategies (e.g., explain a concept, ask a question) and intentions (e.g., motivate the student, get the student to elaborate the answer) to facilitate the feedback generation process. ReMath is a great example of high-quality data collection with human experts towards building better hint generation systems.
TriviaHG Mozafari et al. (2024b) is another hint generation dataset developed by extending the TriviaQA dataset Joshi et al. (2017). Mozafari et al. (2024b) utilize Microsoft’s CoPilot [^2] https://copilot.microsoft.com/ to generate hints due to its retrieval augmented generation (RAG) approach that generated more reliable responses grounded on internet-retrieved documents. TriviaHG includes 10 generated hints for questions that CoPilot is capable on answering. In a follow-up work, Mozafari et al. (2024a) propose HintQA that utilizes these hints as concise context, improving the QA capabilities of LLMs over other context-retrieval and context-generation baselines.
@@
Approaches. For open-ended hint generation of factoid questions, Jatowt et al. (2023) proposed a Wikipedia-based retrieval framework for the “Who?”, “Where?” and “When?” type questions. They propose a popularity-based framework for the "When" question type, where the popularity of an event for the answer year is measured by the count of Wikipedia hyperlinks directing to the event’s website, and a hand-curated template approach for the "Who" and "Where" question types (example hint described in Table 1). On the other hand,
Wang et al. (2023) benchmarked the ReMath dataset by instruction fine-tuning Wei et al. (2021) the language models like Flan-T5 Chung et al. (2022) and GODEL Peng et al. (2022), and using in-context learning Dong et al. (2022) prompts for gpt-3.5-turbo and gpt-4 Achiam et al. (2023). Tack and Piech (2022) and Wang et al. (2023) found the direct use of LLMs to fall short in comparison to human expert responses.
Pal Chowdhury et al. (2024) proposes a hint generation framework for middle-school level math word problems (termed MWPTutor). MWPTutor provides a hint by formulating a question around the next operation to be performed in the state space. This hint is obtained by matching the ongoing response with all possible decomposed solutions obtained by using a language model to decompose the solutions into atomic mathematical operation steps. Current question answering hint generation systems do not personalize the hints to learners’ preferences, learning objectives, or their prior knowledge (Sections 2.1 and 2.2). We discuss how we can improve these hint generation systems to aid the learning process in Section 4.
@@
Evaluation Metrics. All the discussed approaches for hint generation have used human evaluation to assess the quality of the system’s output. Jatowt et al. (2023) conducted a between-subjects study to evaluate their proposed hint generation strategies across different experimental groups. Tack and Piech (2022) proposed the "AI Teacher Test", comparing the generated responses against the teacher responses across three dimensions - “speak like a teacher”, “understand a student”, and “help a student”. They identified that the LLMs are good at conversation uptake (i.e., the first two requirements) but are quantifiably worse than real teachers on several pedagogical dimensions, especially helpfulness to a student. Wang et al. (2023) evaluate the error type identification and feedback response strategy selection as a multi-class classification task and utilize exact match and Cohen’s kappa to measure the accuracy and entropy to measure the output diversity. They also conducted human evaluations for the response generation task and found that all models constrained by knowledge of ground-truth error type and response strategy outperformed their unconstrained counterparts.
For evaluation of hint generation systems in quantitative user studies, Pal Chowdhury et al. (2024) extends the idea of success rate ($S$: fraction of correctly answered questions) and telling rate ($T$: fraction of conversations where the answer was revealed) proposed by Macina et al. (2023). Pal Chowdhury et al. (2024) suggested an adjusted success ($S-T$) to prevent an overtly revealing framework from achieving high performance, and the harmonic mean of success rate and adjusted success rate as the overall tutoring score ($\frac{2S(S-T)}{2S-T}$).
Mozafari et al. (2024b), on the other hand, proposed two learner-agnostic automatic evaluation metrics: convergence to measure the ability of a hint to eliminate wrong candidate answers, and familiarity to measure the recognizability of answer entities. To measure convergence of a hint, they adopt a three-step process - i) generating candidate answers using LLMs, ii) validating the entailment of answer given a hint, and iii) computing an aggregate score for a hint across all candidate answers. For familiarity, they utilize the Page Views of Wikipedia [^3] https://www.wikipedia.org/ articles corresponding to the named entities present in the hint as a measure of global familiarity to the hint normalized across all questions in the TrivialHG corpus Mozafari et al. (2024b).
@@ 4 Roadmap for Future Research in Hint Generation
Great progress has been made in automatic hint generation over the last two decades; however, there is still room for improvement. Existing hint generation frameworks do not personalize the hints to the learner’s prior knowledge (Section 2.1), and are only evaluated in short-term studies. We still do not know the effects of long-term exposure to these interventions on the learners. These frameworks are also limited to certain domains and have not been widely explored in other domains, including different branches of science and social science. To improve upon these factors, we discuss a roadmap for future efforts in automatic hint generation. We propose a computational hint generation framework that draws on education and cognitive sciences (Figure 3).
Figure 3: Roadmap of the proposed computational hint generation framework.
@@ 4.1 Formal Definition
Jatowt et al. (2023) proposed a formal definition for hint generation as follows: Given a question $q$ and its correct answer $a$, the task is to generate a hint $h$, such that $P(a|q,h)-P(a|q)>\epsilon$, where $P(a|q,h)$ denotes the probability of a user answering $q$ if the hint $h$ is given, $P(a|q)$ is the probability of a user answering $q$ without the hint, and $\epsilon$ is a threshold parameter ($\epsilon>0$). This definition emphasizes the hint’s ability to help answer a question and does not incorporate any pedagogical aspects or principles. It also does not take individual preferences into consideration.
Below, we provide a comprehensive definition that draws inspiration from widely adopted cognitive frameworks on human learning, such as Anderson’s adaptive control of thought rational (ACT-R) theory Anderson (2013) and Ausubel’s theory on meaningful learning Ausubel (1963, 1962, 2012). We also incorporate the findings from the qualitative experiments discussed in Section 2, explicitly integrating the alignment of hints to students’ learning objectives and their prior knowledge (Section 2.2), as well as incorporating the pragmatics of a hint (Section 2.1) by accounting for learners’ preferences. We formulate the hint generation task within a tutoring framework with the goal of correctly answering a question. We will later explain how the definition can be extended to generate hints for tasks beyond question answering.
@@
Given a learner $l$ attempting to answer a question $q$, a hint generation system H (Figure 3) generates a hint $h\in\mathcal{H}$ by mapping $\textbf{H}:\mathcal{I}\rightarrow\mathcal{H}$, where =$\mathcal{I}=\{q,a,\mathcal{K}_{q\rightarrow a},D_{q}^{l},\mathcal{L}_{l},\mathcal{F}^{l}_{learning},\mathcal{F}^{l}_{pref}\}$ is the input to the hint generation system having the following elements:
  • $q$: question
  • $a$: correct answer
@@
  • $D_{q}^{l}=\{q,\hat{a}_{1},\hat{h}_{1},\hat{a}_{2},\hat{h}_{2},...,\hat{a}_{k}\}$: ongoing dialogue, where $\hat{a}_{i}$ and $\hat{h}_{i}$ are, respectively, the learner’s past attempts and hints related to $q$,
  • $\mathcal{L}_{l}=\{D_{q_{i}}\}$: learner $l$’s past learning history,
@@
H generates the hint h that does not contain the answer (Equation 1), helps the learner to answer the question (Equation 2), and aligns with the learner’s learning objective(s) (Equation 3).
@@
Both Anderson’s ACT-R theory and Ausubel’s meaningful learning theory imply the significance of connecting the new knowledge required to solve a problem to the existing concepts in the learner’s knowledge base (entailed in Equation 4). We also account for the diversity of learners’ motivation to study and incorporate the notion of improving an individual’s learning objective (Equation 3). The hint generation strategy for a learner aiming to improve their answer accuracy would greatly differ from someone aiming to maximize the diversity of acquired knowledge in a learning session.
Extending the formal definition. The given definition assumes a question-answering setup for the hint generation task, with an additional assumption of the availability of objective answers. We can modify this definition to accommodate for other hint generation settings as described below.
@@
  • For a hint generation system for the writing assistance task, we need to replace the question with the writing task description ($q\rightarrow T$) and ongoing dialogue with an interactive sequence of learner’s writings ($\hat{a_{i}}\rightarrow\hat{w_{i}}$), and past hints ($\hat{h_{i}}$), $\mathcal{D}_{q}^{l}\rightarrow\mathcal{W}_{T}^{l}\mid\mathcal{W}_{T}^{l}\Leftrightarrow\{\hat{w_{1}},\hat{h_{1}},\hat{w_{2}},\hat{h_{2}},...\}$, and the answer with a target rubric for writing evaluation $a\rightarrow\mathcal{R}$.
@@
4.2 Components of an Effective Automatic Hint Generation System